# Effective-medium approach to the resonance distribution of wave scattering in a random point field

### v2

@@ Appendix I Derivation of the diffusion equation
where $\mathcal{V}=\mathcal{B}_{d}(R)$ denotes the spherical volume of the medium.
@@
We first look at the bulk of the medium, that is to say, far away from the boundary ($R-\left\|\boldsymbol{\mathrm{r}}\right\|\gg\ell_{\rm s}$). In this region, it is reasonable to extend the integral to $\mathbb{R}^{d}$, hence neglecting the possible edge effects. We will consider the special case of the boundary equation later. Omitting the constant prefactor $n\sigma$, the convolution integral (189) reads
@@
according to Eq. (71). Since $n\sigma$ is very large, the kernel $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})$ looks like the Dirac delta function $\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime})$, but with a nonzero variance. To exploit this feature, we expand the density $\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle$ in the neighborhood of the observation point $\boldsymbol{\mathrm{r}}$
@@
where we have assumed the implicit summation of repeated indices. We hope that the first three terms of Eq. (192) will suffice to approach $\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle$. The convolution integral (190) then amounts to evaluating the first moments of the distribution $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})$. The zero-order moment reads
@@
is equal to the zero vector, because the distribution is symmetric with respect to the point $\boldsymbol{\mathrm{r}}^{\prime}=\boldsymbol{\mathrm{r}}$. This also means that the particle does not undergo any drift or external forces. In fact, all the odd order moments are zero for the same reason. The second-order moment, which is a $3\times 3$ matrix, reads
@@
Multiplying both sides of Eq. (197) by $(\gamma+n\sigma)$ quickly leads to Eq. (77).
@@
We neglect the higher-order terms in Eq. (200), because we do not want the boundary condition to possess a second order derivative. Indeed, the diffusion equation (197) is of second order in space. Note that, in Eq. (199), the observation point $\boldsymbol{\mathrm{r}}$ is supposed to lie on the boundary, i.e., to satisfy $\boldsymbol{\mathrm{r}}\cdot\boldsymbol{\mathrm{n}}=0$. Without loss of generality, we can set $\boldsymbol{\mathrm{r}}=\boldsymbol{\mathrm{0}}$ due to the translational symmetry of Eq. (199) in the plane of the boundary. The zero-order moment of the kernel reads
@@
which, after some rearrangements, leads to Eq. (84).

### v3

@@ Appendix B Diffusion approximation
where * denotes the convolution product and $\mathcal{V}$ the volume of the medium.
@@
We first look at the bulk of the medium, that is to say, far away from the boundary. In this region, it is reasonable to extend the integral to $\mathbb{R}^{d}$, hence neglecting the possible edge effects. We will consider the special case of the boundary equation later. Omitting the constant prefactor $n\sigma$, the convolution integral (67) reads
@@
according to Eq. (41). Since $n\sigma$ is very large, the kernel $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})$ looks like the Dirac delta function $\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime})$ plus a correction involving the nonzero variance of the kernel (69). To exploit this feature, we may expand the density $\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle$ in the neighborhood of the observation point $\boldsymbol{\mathrm{r}}$
@@
where we have assumed the implicit summation of repeated indices. We suppose that the first three terms of Eq. (70) will suffice to approach $\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle$. The convolution integral (68) then amounts to evaluating the first moments of the distribution $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})$. The zero-order moment reads
@@
is equal to the zero vector, because the distribution is symmetric with respect to the point $\boldsymbol{\mathrm{r}}^{\prime}=\boldsymbol{\mathrm{r}}$. This also means that the particle does not undergo any drift or external forces. In fact, all the odd-order moments are zero for the same reason. The second-order moment, which is a $3\times 3$ matrix, reads
@@
Multiplying both sides of Eq. (75) by $(\gamma+n\sigma)$ readily leads to
$\gamma\left\langle\rho(\boldsymbol{\mathrm{r}},\gamma)\right\rangle=\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}_{0})+\frac{1}{dn\sigma}\nabla^{2}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}},\gamma)\right\rangle\>,$ (77)
whence the first line of Eq. (49) since $\ell_{\rm s}=\tfrac{1}{n\sigma}$.
@@
We neglect the higher-order terms in Eq. (80), because we do not want the boundary condition to possess a second-order derivative. Indeed, the diffusion equation (75) is of second order in space. Note that, in Eq. (79), the observation point $\boldsymbol{\mathrm{r}}$ is supposed to lie on the boundary, i.e., to satisfy $\boldsymbol{\mathrm{r}}\cdot\boldsymbol{\mathrm{n}}=0$. Without loss of generality, we can set $\boldsymbol{\mathrm{r}}=\boldsymbol{\mathrm{0}}$ due to the translational symmetry of Eq. (79) in the plane of the boundary. The zero-order moment of the kernel reads
@@
which, after some rearrangements, leads to
$\left(\gamma+\frac{n\sigma}{2}\right)\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle+\frac{V_{d-1}}{S_{d}}\frac{n\sigma}{\gamma+n\sigma}\boldsymbol{\mathrm{n}}\cdot\boldsymbol{\mathrm{\nabla}}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=0\>.$ (89)
When considering times much longer than the intercollisional time, the variable $\gamma$ in Eq. (89) can be neglected in front of $n\sigma$. In this way, we obtain a version of Eq. (89) which no longer depends on $\gamma$, i.e., giving the time-independent boundary condition
$\frac{n\sigma}{2}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle+\frac{V_{d-1}}{S_{d}}\boldsymbol{\mathrm{n}}\cdot\boldsymbol{\mathrm{\nabla}}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=0\>,$ (90)
on the density $\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle$ for all $\boldsymbol{\mathrm{r}}\in\partial\mathcal{V}$. This is the Robin boundary condition given in the second line of Eq. (49). Since the coefficients of the two terms in Eq. (90) are positive, this condition expresses the fact that the density has an inward-facing gradient. In the far diffusive regime ($n\sigma R\rightarrow\infty$), this condition reduces to a simple Dirichlet boundary condition: $\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=0$. Although this regime is valid for a very large medium, we do not consider this extreme case in this paper.